OpenStudy (jaylelile):

Please help.. I don't even know where to begin with this problem. Two years after an expansion, a company had sales of $420,000. Four years later, sales were $1,220,000. Assuming that sales and time since the expansion are related by a linear equation, determine the expected sales in another 6 years.

OpenStudy (pawanyadav):

Let at time of expansion company has x And which are increased linearly with time.so let they are increased by y per year so the linear equation should be.. x+2y=420,000 And. x+4y=1,220,000 You have two equation and two variable Can you find x,y.

OpenStudy (pawanyadav):

After you find the values of x, y You have to find 1,220,000+6y =? As it is asking for another 6 years after they have reached 1,220,000.

OpenStudy (benlindquist):

its asking how long and ho much

OpenStudy (jaylelile):

so I got x=419,998 and y=200,000 I'm not sure how to put this in a sentence to answer the question.....

OpenStudy (jaylelile):

@mathmale

OpenStudy (jaylelile):

@mathstudent55

OpenStudy (pawanyadav):

You have to find 1,220,000+6y So just plug the value of y

OpenStudy (pawanyadav):

You got wrong values of x,y

OpenStudy (pawanyadav):

Correct value of y=400,000

OpenStudy (pawanyadav):

Got it

OpenStudy (jaylelile):

so I got 3,620,000...

OpenStudy (jaylelile):

@Pawanyadav

OpenStudy (pawanyadav):

Correct